OpenStudy (anonymous):

When surveyed, 5 out of every 8 students liked the food in the cafeteria. If 60 liked the food, how many were surveyed? A. 100 B. 108 C. 98 D. 96

OpenStudy (hoblos):

5/8 = 60/x x=60*8/5 = 96

OpenStudy (anonymous):

thnks!

OpenStudy (hoblos):

yw :)